if one person can complete a task in 2h and another person can complete the same task in 3h, will it take more or less than 2h to complete the task when both people are working?
Found 2 solutions by richwmiller, lwsshak3:
Answer by richwmiller(17219)

it will be a little less than 2 hours.
1/2+1/3=1/x
3/6+2/6=1/x
5/6=1/x
6/5=x
1.2= x


Answer by lwsshak3(11628)

if one person can complete a task in 2h and another person can complete the same task in 3h, will it take more or less than 2h to complete the task when both people are working?
***
sum of individual work rates=work rate working together
1/2+1/3=1/x
3/6++2/6=1/x
5/6=1/x
x=6/5
it will take less than 2h to complete the task when both people are working